OpenStudy (anonymous):

3/4 divided by 5/6 ?

OpenStudy (turingtest):

In general, the rule to know here is\[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times\frac{d}{c}\]in your particular case we have a=3 b=4 c=5 d=6 Plug in the numbers to the proper place in the formula and you get the answer.

OpenStudy (anonymous):

Awesome ! That helps a bunch thank you !

OpenStudy (anonymous):

\[\frac{3/4}{5/6}=\frac{3}{4}*\frac{6}{5}=\frac{9}{10} \]